# “Rosenbluth” separation of the $J/\psi$ near-threshold photoproduction - an access to the gluon Gravitational Form Factors at high $t$

### v2

@@ V Additional remarks and summary
$d\sigma/dt(E_{i},t)\frac{\xi^{2}(E_{i},t)}{F(E_{i})}>d\sigma/dt(E_{j},t)\frac{\xi^{2}(E_{j},t)}{F(E_{j})},\;E_{i}>E_{j}.$ (33)
In particular, as $\xi^{2}(E_{i},t)/F(E_{i})<\xi^{2}(E_{j},t)/F(E_{j}),$222 This inequality follows from Eq.(5) and the kinematical factor $F(E)$ being proportional to $1/(s-m^{2})^{2}$ Guo et al. (2024). it means that:
@@
In the GPD approach our “Rosenbluth” procedure relies on Eq.(1) that comes in Ref. Guo et al. (2024) from some general theoretical assumptions about the $J/\psi$ photoproduction mechanisms, such as gluon exchange and factorization. The observation of the energy independence of the extracted $G$ form factor functions means that the experimental data, within their uncertainties, are consistent with the assumptions in Ref. Guo et al. (2024). In the holographic case, Eq.(2), which relates directly $d\sigma/dt$ to the gGFFs can be interpreted as a more general expression in powers of $\eta$ as in Eq.(14). The energy independence of the extracted $H$ functions supports such an interpretation within the errors of the present data set. We stress that these results are based only on Eqs.(1,14) and no additional information is used. Note that any constraints on the gGFF like fixing parameters on their parametrizations depend on the assumption of a leading-term approximation.
The results from the “Rosenbluth” separation for the parameters of the functions fitted to $G_{0}$, $H_{0}$, and $H_{2}$ (Figs. 4,6) and the results of the global fits in Figs. 2,5 are consistent with each other333 The “Rosenbluth” results for $G_{2}$ are also consistent with the fit parameters in Fig. 2. For the sake of focusing on the analogy between the GPD and holographic approaches, these results are not shown in the paper, as it is the sum $G_{0}+G_{2}$ that has the same relation to the gGFFs as the $H_{2}$ function.. This is a result of using the same data set and the same parametrization for the $G$ and $H$ functions. However, these two methods of extracting the $G$ and $H$ functions are not equivalent. The global fits assume a priori the functional forms used to parametrize the $G$ and $H$ functions. The “Rosenbluth” separation uses only the function fitted to the reference data set. The global fits to all the data used in this case are needed to reduce the systematic errors and, in the GPD case, to extrapolate the reference data to lower $|t|$ values. In the holographic case we can extract the $H$ functions by interpolating the reference cross section without performing a global fit of all the data. The fit of the data points extracted in this way is used to infer (rather than assume) the proper functional form, a squared dipole function, used in the global fits of the reduced cross sections.
@@
In Fig. 7 we have compared the extracted $G$ and $H$ functions with the calculations in the leading-term approximation when using lattice gGFFs. The procedure described in this paper does not use any external constraints and does not depend on the absolute normalization of the corresponding theoretical calculations. The absolute value of the extracted form factor functions comes from the data - more precisely form the slope of the differential cross section as a function of the energy. Therefore we find it remarkable that the data points obtained with two diametric theories are on the same scale as the lattice results. If with higher statistics the agreement between the data and lattice calculations in the overlapping $t$-region is confirmed, this could justify the extraction of the gGFFs at high $t$ directly from the data, complementary to the lattice results available at low $t$.
@@
The results in this paper should not be understood as a proof of the validity of the assumptions in Ref. Guo et al. (2024) and Ref. Mamo and Zahed (2022), rather just as an indication, within the present data uncertainties, of the validity of the scaling as defined by Eqs.(1,14). Such scaling behavior as demonstrated experimentally in this work can be considered as an independent phenomenological observation possibly coming from some more general theoretical assumptions. In any case, the presented results are encouraging and should stimulate the continuation of the experimental studies and further theoretical analysis.
@@ VI ACKNOWLEDGMENTS
We would like to thank Yuxun Guo and Ismail Zahed for the fruitful discussions and explanations of the theories on which the results of this work are based and Simon Taylor for valuable comments.

### v3

@@ V Additional remarks and summary
$d\sigma/dt(E_{i},t)\frac{\xi^{2}(E_{i},t)}{F(E_{i})}$ $>d\sigma/dt(E_{j},t)\frac{\xi^{2}(E_{j},t)}{F(E_{j})},$
$E_{i}>E_{j}.$
In particular, as $\xi^{2}(E_{i},t)/F(E_{i})<\xi^{2}(E_{j},t)/F(E_{j}),$333 This inequality follows from Eq.(5) and the kinematical factor $F(E)$ being proportional to $1/(s-m^{2})^{2}$ Guo et al. (2024). it means that:
@@
In the GPD approach our “Rosenbluth” procedure relies on Eq.(4) that comes in Ref. Guo et al. (2024) from some general theoretical assumptions about the $J/\psi$ photoproduction mechanisms, such as gluon exchange and factorization. The observation of the energy independence of the extracted $G$ form factor functions means that the experimental data, within their uncertainties, are consistent with the assumptions in Ref. Guo et al. (2024). In the holographic case, Eq.(2), which relates directly $d\sigma/dt$ to the gGFFs can be interpreted as a more general expression in powers of $\eta$ as in Eq.(14). The energy independence of the extracted $H$ functions supports such an interpretation within the errors of the present data set. We stress that these results are based only on Eqs.(4,14) and no additional information is used. Note that any constraints on the gGFFs like fixing parameters on their parametrizations depend on the assumption of a leading-term approximation.
The results from the “Rosenbluth” separation for the parameters of the functions fitted to $G_{0}$, $H_{0}$, and $H_{2}$ data points (Figs. 4,6) and the results of the global fits in Figs. 2,5 are consistent with each other444 The “Rosenbluth” results for $G_{2}$ are also consistent with the fit parameters in Fig. 2. For the sake of focusing on the analogy between the GPD and holographic approaches, these results are not shown in the paper, as it is the sum $G_{0}+G_{2}$ that has the same relation to the gGFFs as the $H_{2}$ function.. This is a result of using the same data set and the same parametrization for the $G$ and $H$ functions. However, these two methods of extracting the $G$ and $H$ functions are not equivalent. With the “Rosenbluth” separation we extract data points, while the global fit method assumes a priori the functional forms used to parametrize the $G$ and $H$ functions. The “Rosenbluth” separation uses only the function fitted to the reference data set. The fits to all the data used in this case are needed to reduce the systematic errors and, in the GPD case, to extrapolate the reference data to lower $|t|$ values. In the holographic case we can extract the $H$ functions by interpolating the reference cross section without fitting all the data. The results extracted in this way are similar to those presented above, except for the bigger systematic uncertainties.
@@
In Fig. 7 we have compared the functions fitted to the $G$ and $H$ data points with the calculations in the leading-term approximation when using lattice gGFFs. The procedure described in this paper does not use any external constraints. The absolute value of the extracted form factor functions comes from the slope of the reduced cross sections as a function of the energy. Thus, we find it remarkable that the data points obtained with two diametric theories are on the same scale as the lattice results. If with higher statistics the agreement between the data and lattice calculations in the overlapping $t$-region is confirmed, this could justify the extraction of the gGFFs at high $t$ directly from the data, complementary to the lattice results available at low $t$.
In a recent paper the GPD framework is extended by including Next-to-Leading-Order (NLO) calculations Guo et al. (2025). The authors infer the GFFs for gluons and quarks (the latter enter in the NLO diagrams), parametrized with dipole/tripole functions, by fitting the data and lattice results separately and simultaneously. When fitting the data only, constraints on the gluon/quark GFFs $A_{g,q}(0)$ are applied and the results are found to be generally compatible with the lattice fits. The Jefferson Lab data for $\xi>0.5$ are used in their analysis. The theoretical error at this limit of the skewness parameter is estimated to be $\approx 30\%$. The work presented in our paper can be considered to be an independent and complementary approach that confirms the agreement between the data and lattice results within the present theoretical and experimental uncertainties.
@@
The results in this paper should not be understood as a proof of the validity of the assumptions in Ref. Guo et al. (2024) and Ref. Mamo and Zahed (2022), rather just as an indication, within the present data uncertainties, of the validity of the scaling as defined by Eqs.(4,14). Such scaling behavior as demonstrated experimentally in this work can be considered as an independent phenomenological observation possibly coming from some general expansion in the skewness parameter. In any case, the presented results are encouraging and should stimulate the continuation of the experimental studies and further theoretical analysis.
@@ VI ACKNOWLEDGMENTS
We would like to thank Yuxun Guo, Kiminad Mamo, and Ismail Zahed for the fruitful discussions and explanations of the theories on which the results of this work are based and Simon Taylor for valuable comments. We thank Dimitra Pefkou for providing us with parametrizations of the lattice results.